OpenStudy (anonymous):

y=3sin(2)(theta−pie/6 )−1/2 -Can someone pleas show me how to graph this?

OpenStudy (anonymous):

\[3\sin[2(\theta - \frac{\pi}{6})] - \frac{1}{2}\]?